using matrices to solve the systems.
x+5y= 0
x+6y+z= 1
5x-y-z+= -101
this is what i came up with but it doesn't look right, plz help
1 5 0 0
0 1 1 1
0 0 1 127/27

Answer by AnlytcPhil(1808)

x+5y= 0
x+6y+z= 1
5x-y-z+= -101
Note: the number I write on the left
of each matrix tells what I'm going to 
multiply a row by and add it to the
row with a 1 left of it to get the
next matrix:

-1[1  5  0 |    0]
 1[1  6  1 |    1]
  [5 -1 -1 | -101]

-5[1  5  0 |    0]
  [0  1  1 |    1]
 1[5 -1 -1 | -101]

 1[1   5  0 |    0]
-5[0   1  1 |    1]
  [0 -26 -1 | -101]

  [1   0 -5 |   -5]
26[0   1  1 |    1]
 1[0 -26 -1 | -101]

   [1   0 -5 |  -5]
   [0   1  1 |   1]
÷25[0   0 25 | -75]

  1[1   0 -5 | -5]
   [0   1  1 |  1]
  5[0   0  1 | -3] 

   [1   0  0 |-20]
  1[0   1  1 |  1]
 -1[0   0  1 | -3]

   [1   0  0 |-20]
   [0   1  0 |  4]
   [0   0  1 | -3]

x=-20, y=4, z=-3

Edwin